# Towards Motion Compensation in Autonomous Robotic Subretinal Injections

### v1

@@ Abstract
Exudative (wet) age-related macular degeneration (AMD) is a leading cause of vision loss in older adults, typically treated with intravitreal injections. Emerging therapies, such as subretinal injections of stem cells, gene therapy, small molecules or RPE cells require precise delivery to avoid damaging delicate retinal structures. Autonomous robotic systems can potentially offer the necessary precision for these procedures. This paper presents a novel approach for motion compensation in robotic subretinal injections, utilizing real-time Optical Coherence Tomography (OCT). The proposed method leverages B5-scans, a rapid acquisition of small-volume OCT data, for dynamic tracking of retinal motion along the Z-axis, compensating for physiological movements such as breathing and heartbeat. Validation experiments on ex vivo porcine eyes revealed challenges in maintaining a consistent tool-to-retina distance, with deviations of up to 200 $\mu m$ for 100 $\mu m$ amplitude motions and over 80 $\mu m$ for 25 $\mu m$ amplitude motions over one minute. Subretinal injections faced additional difficulties, with horizontal shifts causing the needle to move off-target and inject into the vitreous. These results highlight the need for improved motion prediction and horizontal stability to enhance the accuracy and safety of robotic subretinal procedures.
@@ I INTRODUCTION
Figure 1: OCT B-scan images showing the simulated breathing motion transmitted to the retina with the needle following the same motion while trying to keep a constant distance to it. Red and green lines are the starting position of the needle tip and retina respectively. As our motion compensation is based on comparing the current with the previously acquired volumes, we see a latency in the needle’s motion (scans 3 and 5). The needle continues to move in the direction opposite the retina. The pink and cyan lines are the internal limiting membrane (ILM) and retinal pigment epithelium (RPE) layers of the retina respectively.
@@
Currently, the primary treatment for wet AMD involves injecting anti-VEGF (vascular endothelial growth factor) drugs directly into the vitreous cavity [2] This symptomatic treatment delays disease progression but frequent intravitreal injections are costly and inconvenient to the patients and have serious risks, such as endophthalmitis [3].
@@
Freehand subretinal injection remains a very challenging task. This is largely due to human physiological limits, including hand tremor, which is about 182 $\mu m$ [8], making precise needle tip positioning extremely difficult inside the, on average 250 $\mu m$ thick, retina. Another major challenge is the lack of proper needle tip visualization relative to the target area [5].
@@
To filter out the human hand tremor, several surgical robotic platforms have been developed including the Steady Hand Eye Robot (SHER) [9], Intraocular Robotic Interventional and Surgical System (IRISS) [10], Preceyes Surgical System [11], to name a few [12, 13, 14]. The SHER is a 5-degree-of-freedom (DoF) surgical robotic system developed at Johns Hopkins University. Several versions of these robots are built, such as SHER 2.0 and SHER 2.1, which are serial manipulators and SHER 3.0, which has a compact parallel mechanism [15]. These robots are typically controlled in a cooperative strategy with an admittance-based control algorithm in which the force applied by the user’s hand to the robot handle is measured by a force/torque sensor and interpreted as the desired end-effector velocity of the robot [16]. Cooperative control strategy involves direct visualization of the surgical site through a microscope followed by manual guidance of the robotic tool, which is a challenging task due to constrained visualization in a submillimeter workspace. To resolve this issue several works are developed to autonomously navigate the robotic system to the target area using deep learning methods integrated with intraoperative optical coherence tomography (iOCT) imaging, which can provide high-resolution cross-sectional 2D B-scans or volumetric 3D C-scans of the retinal layers, enabling submillimeter precision for robotic tool placement and control [17, 18, 19, 20].
@@
Although these methods demonstrate promising outcomes in guiding the cannula along a pre-planned path to a target site in ex-vivo porcine eye experiments, they do not account for the dynamic tissue deformation during the insertion process. Tool-tissue interaction during subretinal injection causes deformation in the retina, dynamically shifting the target point. Hence, a pre-planned trajectory with a fixed target point is not a realistic assumption and cannot precisely guide the cannula tip to a desired target. Of note, shifts in the target point along the $X$ and $Y$ axes are less concerning, but precise control over the $Z$-axis (perpendicular to the retina layer) remains essential.
@@
In our prior work, we addressed the retinal tissue deformation along the $Z$-axis by real-time feedback of the tissue deformation and the target point movement, using our novel OCT-based tracking method and studied its feasibility in an ex-vivo porcine eye experiment [21]. This depth accuracy ensures that injections consistently reach the correct tissue layer, especially near the RPE, where maintaining a specific depth is critical to delivering effective treatment without risking tissue damage. This approach represents an initial advancement in robotic subretinal injection adaptable to retina deformation; by creating and regularly updating a virtual target layer that is defined at a relative depth between the ILM and RPE layers, it enhances procedural safety and ensures that the fragile RPE cells remains intact during the insertion. However, the porcine eyes were placed on a 3D-printed stand and did not move during the experiment. In the actual human eye, retina has alternating movement, caused predominantly by breathing and the heartbeat. This motion can be a source of unintended damage to the delicate retinal cells if not accounted for. For example, the retinal movement along the axial direction ($Z$-axis, perpendicular to the retina layer) caused by heartbeat and breathing is reported to be of the amplitude of 81$\pm$3.5 $\mu m$ with a frequency of approximately 1 $Hz$. Additionally, the amplitude of the retina’s axial motion is recorded as 21.3$\pm$8 $\mu m$ while the subject is lying down to minimize the effects of head movements [22].
@@
This work presents a novel approach for motion compensation in autonomous robotic eye surgery, explicitly targeting subretinal injections. Our method leverages real-time acquisition and segmentation of small OCT volumes, called B5-scans, to track retinal movement along the $Z$-axis. This allows us to dynamically adjust the position of the SHER robotic system’s end-effector, maintaining the position of the surgical tools in relation to the retina (Fig. 1). To our knowledge, this is the first implementation of an OCT-based motion compensation technique in the context of autonomous robotic surgery. However, it is important to note that this approach is an initial proof-of-concept. While our method successfully demonstrates the feasibility of real-time motion compensation, it remains a preliminary step towards more advanced and reliable techniques for automated subretinal interventions. The contributions of this work consist of:
@@
  • The first implementation of an autonomous motion compensation system for robotic eye surgery, specifically targeting subretinal injections. This system uses real-time OCT volume acquisition and segmentation, facilitated by small-volume scans termed B5-scans. By tracking the motion of retinal layers in real-time, our method enables precise control of the robotic end-effector to match retinal movements.
@@ II METHODS
Our motion compensation is based on the real-time processing of OCT volumes. As depicted in Fig. 2, the incoming OCT volumes are segmented and point clouds of the ILM, RPE, and needle surfaces are generated from the segmentation results. By comparing the median $Z$-axis position of the layer points to the previous point cloud the direction of the retina motion can be inferred. A velocity is applied to the robot end-effector in the positive or negative $Z$ direction in accordance with the retinal motion, allowing the needle to be kept at the same position relative to the retina.
@@ II-A B5-scans
Since OCT microscope systems with real-time C-scan acquisition capabilities are not yet commercially available, we use small volume scans, which we call B5-scans, instead of large C-scans for imaging the surgical area. By combining five equally spaced B-scans of a 0.1 $\times$ 4 mm scan area, we can create a B5-scan. Because of the small number of B-scans, a B5-scan can be acquired in approximately 0.1 seconds.
For comparison, using the Leica Proveo 8 with the EnFocus OCT imaging system acquiring a conventional large C-scan (4 $\times$ 4 mm scan area, 200 $\times$ 1000 $\times$ 1024 pixel resolution) takes 6 seconds. The fast acquisition times provides real-time information necessary for robotic applications.
@@ II-B OCT Segmentation Network
To calculate retina’s motion, the positions of the retinal layers need to be known. We used an OCT B-scan segmentation network previously introduced in [21], where we described our training data and procedure. This network segments the top layer surface of the ILM, RPE, and the needle. In this work we present the network parameters, in Table I, which were not included in our previous work.
@@ II-C Surface Point Cloud Generation
As this work is an addition to the autonomous robotic subretinal injection pipeline in [21], we chose to use the surface point cloud generated for calculating the layer depth. This reduces the number of points significantly while potentially removing mislabeled pixels.
As we are only interested in the topmost surfaces of the needle, ILM and RPE layers, we create a point cloud by taking the first occurrence of each class along the vertical axis. A more detailed explanation of this procedure can be found in [21].
@@ II-D Retina Motion Compensation
To calculate the motion of the retina in the $Z$ direction, we use the median $Z$-axis position of the ILM layer points from the point cloud generated in the previous step. The choice of the median, rather than the mean, as a central measure is deliberate: it minimizes the impact of outliers that could lead to inaccurate conclusions about retinal movement. By focusing on the median $Z$-axis position, we ensure a more robust representation of the ILM’s position, despite potential noise or extreme values in the data.
To estimate the direction of the retina’s motion, we compare this median $Z$-axis position to the value from the previous B5-scan. If the current median $Z$-axis position is greater than the previous one, the retina is moving downward along the $Z$-axis. Conversely, if the current value is smaller than the previous, it suggests an upward movement along the $Z$-axis.
Based on the direction of the retinal motion, we apply a constant velocity in the same direction as the robot end-effector along the $Z$-axis.
Figure 3: The experimental setup includes, (a) the SHER 2.0, an iOCT microscope, a piezo-actuated linear stage to generate a sinusoidal motion along the $Z$ direction (up and down) simulating effects of human breathing on the eye, a syringe pump, a syringe with a 42 gauge needle, and (b) a closer view of the open sky porcine eye in the 3D printed holder.
Figure 4: The preparation steps of the porcine eye for the experiment. The unprocessed ex vivo porcine eye (a) is cleaned by excision of the surrounding extraocular muscles (b), then bisected to remove the anterior segment including cornea, pupil, and lens, thereby exposing the retina and vitreous (c) and placed in the 3D printed holder attached to the piezo-actuated linear stage (d).
@@ III-A Experimental Setup
Fig. 3 presents the experimental setup, which includes the SHER 2.0 whose joint-level velocities are controlled by a Galil motion controller (Galil 4088, Galil, Rocklin, CA, USA), an iOCT surgical microscope (Proveo 8, Leica Microsystems, Germany), a syringe pump (PHD2000, Harvard Apparatus, USA), a piezo-actuated linear stage (Q-Motion Stages, PI Motion and Positioning, MA, USA), a syringe with a 42 gauge needle (INCYTO Co., South Korea), a 3D printed eye holder, and porcine eyes. The robot velocity controller is implemented in C++ using the CISST-SAW libraries [23] and the deep learning algorithms are implemented in Python using PyTorch and MONAI frameworks. All these components communicate via the Robot Operating System (ROS) through a TCP-IP connection.
@@ III-B Experimental Procedure
In our experiments, we used ex-vivo open-sky porcine eyes. The preparation steps of these eyes can be seen in Fig. 4.
@@
For our initial experiments, we tested amplitudes of 25, 50, and 100 $\mu m$ to replicate varying degrees of retinal motion. In these preliminary tests, the needle was positioned above the retina, without contact, and we aimed to maintain a constant distance between the needle and the retinal surface.
@@
(a) 100 $\mu m$ amplitude retina motion, 0.8 mm/s robot velocity
(b) 50 $\mu m$ amplitude retina motion, 0.4 mm/s robot velocity
(c) 25 $\mu m$ amplitude retina motion, 0.2 mm/s robot velocity
(d) 25 $\mu m$ amplitude retina motion, 0.3 mm/s robot velocity
@@
Based on this formula the velocity of the 25, 50, 100 $\mu m$ amplitude motions are 0.2, 0.4 and 0.8 mm/s respectively. Although the 0.2 mm/s velocity is theoretically correct for a 25 $\mu m$ amplitude motion, we also tested applying a slightly higher velocity of 0.3 mm/s to the robot end-effector.
In the second experiment, we set the eye motion amplitude to 100 $\mu m$ and positioned the needle above the retina. Using the method outlined in our previous paper [21], we performed needle insertions and activated the breathing compensation described here once the target insertion depth was achieved. We moved the robot end-effector with a 0.8 mm/s velocity based on the results of our previous experiment. Following insertion, a 0.1 ml fluid injection was administered using a syringe pump with an infusion rate of 1 ml/min to evaluate the insertion’s quality. At the same time, both the eye and needle were in motion. This procedure was repeated for seven separate injections. We considered the injection a success if the injected water flowed under the retina causing retinal detachment and retinal bleb formation, examples of successful injections can be seen in Fig. 7.
@@ IV RESULTS
In Fig. 5, we present the $Z$-axis positions of the linear stage controller and needle over time during our first experiment, where we tried to maintain a constant distance between the needle tip and the retina for 60 seconds. Both values are zeroed at their starting positions to allow direct visual comparison. In Fig. 5 (c) and (d), we show results from the 25 $\mu m$ amplitude motion experiments with both 0.2 and 0.3 mm/s robot velocities, respectively.
For the 100 $\mu m$ amplitude motion, we see approximately a 200 $\mu m$ vertical deviation that starts around the 40-second mark. The 50 $\mu m$ amplitude motion performs better at following the linear stages motion.
In Fig. 5 (c) we can see that the theoretically correct 0.2 mm/s end-effector velocity does not have a similar amplitude to the actual linear stage motion. Although using 0.3 mm/s end-effector velocity, we can match the amplitude of the reference motion better, we also have high vertical shift in this experiment, with a maximum of 80 $\mu m$.
@@
Figure 7: Examples of the retinal detachment caused by the bleb (yellow) creation after succesful subretinal injection. The needle is shown in red.

### v2

@@ Abstract
Exudative (wet) age-related macular degeneration (AMD) is a leading cause of vision loss in older adults, typically treated with intravitreal injections. Emerging therapies, such as subretinal injections of stem cells, gene therapy, small molecules and RPE cells require precise delivery to avoid damaging delicate retinal structures. Robotic systems can potentially offer the necessary precision for these procedures. This paper presents a novel approach for motion compensation in robotic subretinal injections, utilizing real time Optical Coherence Tomography (OCT). The proposed method leverages B5-scans, a rapid acquisition of small-volume OCT data, for dynamic tracking of retinal motion along the Z-axis, compensating for physiological movements such as breathing and heartbeat. Validation experiments on ex vivo porcine eyes revealed challenges in maintaining a consistent tool-to-retina distance, with deviations of up to 200 µm for 100 µm amplitude motions and over 80 µm for 25 µm amplitude motions over one minute. Subretinal injections faced additional difficulties, with phase shifts causing the needle to move off-target and inject into the vitreous. These results highlight the need for improved motion prediction and horizontal stability to enhance the accuracy and safety of robotic subretinal procedures.
@@ I INTRODUCTION
Figure 1: OCT B-scan images showing the simulated breathing motion transmitted to the retina with the needle following the same motion while trying to keep a constant distance from it. Red and green lines represent the starting positions of the needle tip and retina respectively. Our motion compensation is based on comparing the current with the previously acquired volumes, so we see a latency in the needle’s motion (scans 3 and 5). The needle continues to move in the direction opposite the retina. The pink and cyan lines are the internal limiting membrane (ILM) and retinal pigment epithelium (RPE) layers of the retina respectively.
@@
Currently, the primary treatment for wet AMD involves injecting anti-VEGF (vascular endothelial growth factor) drugs directly into the vitreous cavity [7]. This symptomatic treatment delays disease progression but frequent intravitreal injections are costly and inconvenient to the patients and have serious risks, such as endophthalmitis [8].
@@
Freehand subretinal injection remains challenging due to human physiological limits, including hand tremor, which is about 182 µm [13], making precise needle tip positioning extremely difficult inside the, on average 250 µm thick, retina. Another major challenge is the lack of proper needle tip visualization relative to the target area [10].
@@
Several surgical robotic platforms have been developed, including the Steady Hand Eye Robot (SHER) [14], Intraocular Robotic Interventional and Surgical System (IRISS) [15], Preceyes Surgical System [16], to name a few [17, 18, 19]. All these platforms filter out physiological hand tremor. The SHER is a 5-degree-of-freedom (DoF) surgical robotic system developed at Johns Hopkins University. Several versions of these robots have been built, such as SHER 2.0 and SHER 2.1, which are serial manipulators and SHER 3.0, which has a compact parallel mechanism [20]. These robots are typically controlled by a cooperative strategy with an admittance-based control algorithm in which the force applied by the user’s hand to the robot handle is measured by a force/torque sensor and interpreted as the desired end-effector velocity of the robot [21, 22]. Cooperative control strategy involves direct visualization of the surgical site through a microscope followed by manual guidance of the robotic tool, which is a challenging task due to constrained visualization in a submillimeter workspace. Prior work has been performed to autonomously navigate the robotic system to the target area using deep learning methods integrated with intraoperative optical coherence tomography (iOCT) imaging, which provides high-resolution cross-sectional 2D B-scans or volumetric 3D C-scans of the retinal layers, enabling submillimeter precision for robotic tool placement and control [23, 24, 25, 26].
@@
Although these methods demonstrate promising outcomes in guiding the cannula along a pre-planned path to a target site in ex vivo porcine eye experiments, they do not account for dynamic tissue deformation during insertion. Tool-tissue interaction during subretinal injection causes deformation in the retina, dynamically shifting the target point. Hence, a pre-planned trajectory with a fixed target point is not a realistic assumption and cannot precisely guide the cannula tip to a desired target. Of note, shifts in the target point along the $X$ and $Y$ axes are less concerning, but precise control over the $Z$-axis (perpendicular to the retina layer) remains essential.
@@
Our prior work, addressed the retinal tissue deformation along the $Z$-axis by real time feedback of the tissue deformation and the target point movement. We used our novel OCT-based tracking method and studied its feasibility in an ex vivo porcine eye experiment [27]. This depth accuracy ensures that injections consistently reach the correct tissue layer, especially near the RPE, where maintaining a specific depth is critical to delivering effective treatment without risking tissue damage. This approach represents an initial advancement in robotic subretinal injection adaptable to retina deformation; by creating and regularly updating a virtual target layer that is defined at a relative depth between the ILM and RPE layers, it enhances procedural safety and ensures that the fragile RPE cells remain intact during the insertion. We placed the porcine eyes on a 3D-printed stand and they did not move during the experiment. In the human eye, the retina has alternating movement, caused predominantly by breathing and the heartbeat. This motion can be a source of unintended damage to the delicate retinal cells if not accounted for. For example, the retinal movement along the axial direction ($Z$-axis, perpendicular to the retina layer) caused by heartbeat and breathing is reported to be of the amplitude of 81$\pm$3.5 µm with a frequency of approximately 1 Hz. Additionally, the amplitude of the retina’s axial motion is recorded as 21.3$\pm$8 µm while the subject is lying down to minimize the effects of head movements [28].
@@
This work presents a novel approach for motion compensation in autonomous robotic eye surgery, explicitly targeting subretinal injections. Our method leverages real time acquisition and segmentation of small OCT volumes, called B5-scans, to track retinal movement along the $Z$-axis. This allows us to dynamically adjust the position of the SHER robotic system’s end-effector, maintaining the position of the surgical tools in relation to the retina (Fig. 1). To our knowledge, this is the first implementation of an OCT-based motion compensation technique in the context of autonomous robotic surgery. However, it is essential to note that this approach is an initial proof-of-concept. While our method successfully demonstrates the feasibility of real time motion compensation, it remains a first step towards more advanced and reliable techniques for automated subretinal interventions. The contributions of this work consist of:
@@
  • The first implementation of an autonomous motion compensation system for robotic eye surgery, specifically targeting subretinal injections. This system uses real time OCT volume acquisition and segmentation, facilitated by small-volume scans termed B5-scans. By tracking the motion of retinal layers in real time, our method enables precise control of the robotic end-effector to match retinal movements.
@@ II METHODS
Our motion compensation is based on the real time processing of OCT volumes. As depicted in Fig. 2, the incoming OCT volumes are segmented and point clouds of the ILM, RPE, and needle surfaces are generated from the segmentation results. The direction of the retina’s motion is determined by comparing the median $Z$-axis position of the layer points to the previous point cloud. A velocity is applied to the robot end-effector in the positive or negative $Z$ direction in accordance with the retinal motion, allowing the needle to be maintained at the same relative position to the retina.
@@ II-A B5-scans
Since OCT microscope systems with real time C-scan acquisition capabilities are not yet commercially available, we use small volume scans, which we call B5-scans, instead of large C-scans for imaging the surgical area. By combining five equally spaced B-scans of a 0.1 $\times$ 4 mm scan area, we can create a B5-scan. Because of the small number of B-scans, B5-scans are acquired in approximately 0.1 seconds.
For comparison, using the Leica Proveo 8 with the EnFocus OCT imaging system acquiring a conventional large C-scan (4 $\times$ 4 mm scan area, 200 $\times$ 1000 $\times$ 1024 pixel resolution) takes 6 seconds. The fast acquisition times provide real time information necessary for robotic applications.
@@ II-B OCT Segmentation Network
The positions of the retinal layers need to be known to calculate retinal motion. We used an OCT B-scan segmentation network previously introduced in [27], where we described our training data and procedure. This network segments the top layer surface of the ILM, RPE, and the needle. This work presents the network parameters in Table I, not included in our previous work.
@@ II-C Surface Point Cloud Generation
As this work is an addition to the autonomous robotic subretinal injection pipeline in [27], we chose to use the generated surface point cloud for calculating the layer depth. This reduces the number of points significantly while potentially removing mislabeled pixels.
As we are only interested in the topmost surfaces of the needle, ILM, and RPE layers, we create a point cloud by taking the first occurrence of each class along the vertical axis. A more detailed explanation of this procedure can be found in [27].
@@ II-D Retina Motion Compensation
To calculate the motion of the retina in the $Z$ direction, we use the median $Z$-axis position of the ILM layer points from the point cloud generated in the previous step. The choice of the median, rather than the mean, is deliberate: it minimizes the impact of outliers that could lead to inaccurate conclusions. Focusing on the median $Z$-axis position ensures a more robust representation of the ILM’s position, despite potential noise or extreme values in the data.
To estimate the direction of the retina’s motion, we compare this median $Z$-axis position to the value from the previous B5-scan. Suppose the current median $Z$-axis position is greater than the previous one. In that case, the retina is moving downward along the $Z$-axis and the opposite is true if the $Z$-axis position is smaller than the last. Comparing the positions of the incoming scans with the previous one, instead of the first scan, overcomes the limitations of intentional or unintentional microscope motion, for example by adjusting the scan area for better visualization.
Based on the direction of the retinal motion, we apply a constant velocity in the same direction as the robot end-effector along the $Z$-axis. The applied velocity is updated approximately every 0.11 seconds after the acquisition and processing of B5-scans.
Figure 3: The experimental setup includes, (a) the SHER 2.0, an OCT microscope, a piezo-actuated linear stage to generate a sinusoidal motion along the $Z$ direction (up and down) simulating effects of human breathing on the eye, a syringe pump, a syringe with a 42 gauge needle, and (b) a closer view of the open sky porcine eye in the 3D printed holder.
Figure 4: Examples of the retinal detachment caused by the bleb (yellow) creation after successful subretinal injection. The needle is shown in red.
@@ III-A Experimental Setup
Fig. 3 presents the experimental setup, which includes the SHER 2.0 whose joint-level velocities are controlled by a Galil motion controller (Galil 4088, Galil, Rocklin, CA, USA), an OCT surgical microscope with a 5 mm imaging depth in tissue (Proveo 8, Leica Microsystems, Germany), a syringe pump (PHD2000, Harvard Apparatus, USA), a piezo-actuated linear stage (Q-Motion Stages, PI Motion and Positioning, MA, USA), a syringe with a 42 gauge needle (INCYTO Co., South Korea), a 3D printed eye holder, and porcine eyes. The robot velocity controller is implemented in C++ using the CISST-SAW libraries [29], and the deep learning algorithms are implemented in Python using PyTorch and MONAI [30] frameworks. All these components communicate via the Robot Operating System (ROS) through a TCP-IP connection.
@@ III-B Experimental Procedure
For our initial experiments, we tested amplitudes of 25, 50, and 100 µm to replicate varying degrees of retinal motion. In these preliminary tests, the needle was positioned at a random point above the retina, without contact, and we aimed to maintain the same distance between the needle and the retinal surface as in the starting position.
@@
(a) 100 µm amplitude retina motion, 0.8 mm/s robot velocity
(b) 50 µm amplitude retina motion, 0.4 mm/s robot velocity
(c) 25 µm amplitude retina motion, 0.2 mm/s robot velocity
(d) 25 µm amplitude retina motion, 0.3 mm/s robot velocity
@@
Based on this formula the velocity of the 25, 50, 100 µm amplitude motions are 0.2, 0.4 and 0.8 mm/s respectively. Although the 0.2 mm/s velocity is theoretically correct for a 25 µm amplitude motion, we also tested applying a slightly higher velocity of 0.3 mm/s to the robot end-effector.
In the second experiment, we set the eye motion amplitude to 100 µm and positioned the needle above the retina. Using the method outlined in our previous paper [27], we performed needle insertions and activated the breathing compensation described here once the target insertion depth was achieved. We moved the robot end-effector with a 0.8 mm/s velocity based on the results of our previous experiment. Following insertion, a 0.1 ml fluid injection was administered using a syringe pump with an infusion rate of 1 ml/min to evaluate the insertion’s quality. At the same time, both the eye and needle were in motion. This procedure was repeated for seven separate injections. We considered the injection successful if the injected water flowed under the retina causing retinal detachment and retinal bleb formation. Examples of successful injections can be seen in Fig. 4.
@@ IV RESULTS
In Fig. 5, we present the $Z$-axis positions of the linear stage controller and needle over time during our first experiment, where we tried to maintain a constant distance between the needle tip and the retina for 60 seconds. Both values are zeroed at their starting positions to allow direct visual comparison. In Fig. 5 (c) and (d), we show results from the 25 µm amplitude motion experiments with 0.2 and 0.3 mm/s robot velocities, respectively.
For the 100 µm amplitude motion, we see approximately a 200 µm vertical deviation that starts around the 40-second mark. The 50 µm amplitude motion performs better at following the linear stages motion.
In Fig. 5 (c) we can see that the theoretically correct 0.2 mm/s end-effector velocity does not have a similar amplitude to the actual linear stage motion. Although using 0.3 mm/s end-effector velocity, we can better match the amplitude of the reference motion. This experiment also has ahigh vertical shift, with a maximum of 80 µm.